OpenStudy (luigi0210):

Differentiate:

OpenStudy (luigi0210):

\[\LARGE y=\frac{4x^3-7x}{5x^2+2}\]

OpenStudy (anonymous):

quotient rule

OpenStudy (luigi0210):

If I'm not mistaken you use quotient rule right? So: \[\LARGE y'=\frac{(5x^2+2)(4x^3-7x)'-(5x^2+2)'(4x^3-7x)}{(5x^2+2)^2}\]

OpenStudy (anonymous):

\[\huge\frac{ f(x)g'(x)-g(x)f'(x) }{f(x)^2}\]

OpenStudy (freethinker):

you can also use product rule, which implements another rule

OpenStudy (anonymous):

implicit differentiation dy/dx y'=20x^4++59x^2-14/(5x^2+2)^2

OpenStudy (freethinker):

y' = 4x^3 - 7x * (5x^2 + 2)^(-1)

OpenStudy (freethinker):

whichever you are most comfortable with you can even differentiate using the definition

OpenStudy (shamil98):

u already know how to do dis

OpenStudy (luigi0210):

Thanks guys @shamil98 Shh xD

OpenStudy (luigi0210):

I'm curious as to the implicit differentiation approach tho.

OpenStudy (freethinker):

just like the other rules or ways of solving it, it would not be optimal to do so besides, implicit differentiation just means you are waiting for the value of the unknown, which in this case is the x y = x^2 y' = 2x * x'

OpenStudy (anonymous):

\(5x^2y+2y-4x^3+7x=0\) now use implicit

OpenStudy (luigi0210):

Oh, now I see.